If In=int_0 to ♾️ {(e^-x)(sin^n) (x) dx } , prove that (1+n^2)In=n(n-1)In-2 for n≥2 .

1
Expert's answer
2022-03-18T09:25:57-0400

Solution

"In = \int\limits_0^\infty {{e^{ - x}}\sin \left( {nx} \right)\,dx} \"


We solve the given integral by integration by parts method


"u = {e^{ - x}}\" and "dv = \sin \left( {nx} \right)\,dx\"


"du = - {e^{ - x}}dx\" and "v = - \frac{{\cos \left( {nx} \right)}}{n}"


Then 


"In = \int\limits_0^\infty {{e^{ - x}}\sin \left( {nx} \right)\,dx} \"


"In = \left[ { - \frac{{{e^{ - x}}\cos \left( {nx} \right)}}{n}} \right]_0^\infty - \int\limits_0^\infty {\left( { - \frac{{\cos \left( {nx} \right)}}{n}} \right)\left( { - {e^{ - x}}} \right)\,dx}"


"In = - \frac{1}{n}\left[ {\frac{{\cos \left( {nx} \right)}}{{{e^x}}}} \right]_0^\infty - \frac{1}{n}\int\limits_0^\infty {{e^{ - x}}\cos \left( {nx} \right)\,dx}"


"In = - \frac{1}{n}\left[ {\frac{{\cos \left( {n\left( \infty \right)} \right)}}{{{e^\infty }}} - \frac{{\cos \left( {n\left( 0 \right)} \right)}}{{{e^{\left( 0 \right)}}}}} \right] - \frac{1}{n}\int\limits_0^\infty {{e^{ - x}}\cos \left( {nx} \right)\,dx}"


"In = - \frac{1}{n}\left[ {0 - 1} \right] - \frac{1}{n}\int\limits_0^\infty {{e^{ - x}}\cos \left( {nx} \right)\,dx}"


"In = \frac{1}{n} - \frac{1}{n}\int\limits_0^\infty {{e^{ - x}}\cos \left( {nx} \right)\,dx}"


 "\begin{array}{l}
In = \frac{1}{n} - \frac{1}{n}\int\limits_0^\infty {{e^{ - x}}\cos \left( {nx} \right)\,dx} \\
\end{array}\"


 "\begin{array}{l}
In = \frac{1}{n} - \frac{1}{n}\left( I \right) & & & ... & \left( 1 \right)
\end{array}\"


Now consider


"I = \int\limits_0^\infty {{e^{ - x}}\cos \left( {nx} \right)\,dx} \"


Solving by integration by parts 


"u = {e^{ - x}}" and  "dv = \cos \left( {nx} \right)\,dx"


"du = - {e^{ - x}}dx\" and  "v = \frac{{\sin \left( {nx} \right)}}{n}"


Then


"I = \left[ {\frac{{{e^{ - x}}\sin \left( {nx} \right)}}{n}} \right]_0^\infty - \int\limits_0^\infty {\frac{{ - {e^{ - x}}\sin \left( {nx} \right)}}{n}dx}"


"I = \frac{1}{n}\left[ {\frac{{\sin \left( {nx} \right)}}{{{e^x}}}} \right]_0^\infty + \frac{1}{n}\int\limits_0^\infty {{e^{ - x}}\sin \left( {nx} \right)dx}"


"I = \frac{1}{n}\left[ {0 - 0} \right]_0^\infty + \frac{1}{n}\left( {In} \right)"


"I = \frac{1}{n}\left( {In} \right)"



Hence from (1), we have 


"In = \frac{1}{n} - \frac{1}{n}\left( {\frac{1}{n}\left( {In} \right)} \right)\\"


"In = \frac{1}{n} - \frac{1}{{{n^2}}}\left( {In} \right)\\"


"In\left( {1 + \frac{1}{{{n^2}}}} \right) = \frac{1}{n}\\"


"In = \frac{1}{n} \times \frac{{{n^2}}}{{1 + {n^2}}}\\"


"In = \frac{n}{{1 + {n^2}}}"


Therefore,


"\int\limits_0^\infty {{e^{ - x}}\sin \left( {nx} \right)\,dx} 
 = \frac{n}{{1 + {n^2}}}"